3. Simplify : 1 7 (i) 22/3 $ 21/5 (ii) b l 33 (iii) 111/2 (iv) 1/2 $ 1/2 7 8 111/4

3. Simplify :
1 7
(i) 22/3 $ 21/5 (ii) b l
33
(iii) 111/2 (iv) 1/2 $ 1/2
7 8
111/4

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) 22/3 $ 21/5 = 2 2/3 +1/5 = 213/15 (ii) b 313 l7 = (313)7 = 331#7 = 3121 = 3−21 (iii) 11111/1/24 = 111/2 −1/4 = 111/4 (iv) 71/2 $ 81/2 = (7.8)1/2 = 561/2